### v1

@@ 1.1 Our Contributions
We study a stochastic combinatorial bandit problem over $T$ rounds with $K$ arms and $U$ users. Each arm $k$ has a cost $c_{k}$, and the learner can select any subset of arms within a budget $B$, and then assign users to the selected arms. Each user derives a stochastic satisfaction from each arm, and the goal is to maximize total user satisfaction. As the general case is NP-complete, we focus on SP preferences (see LABEL:{def:single-peaked-def}), which enable efficient solutions. Our main contributions are as follows.
@@
We consider the case of known SP order and user peaks (but not the cardinal preferences). Such an assumption is justified if, for example, preferences rely on a known ideological spectrum where user peaks correspond to stated or inferred inclinations. We design an efficient UCB-based algorithm, termed MvM, that achieves $\tilde{O}(U\sqrt{TK})$ regret. The key property used in our approach is the existence of a maximal preferences matrix in the confidence set of all plausible SP matrices—that is, a matrix that element-wise dominates all others in the set. We use this maximal matrix as the basis for optimistic matching using our offline algorithm.
@@
Lower bounds
To demonstrate that SP preferences do not trivialize the learning problem, we establish several lower bounds on the worst-case regret. We show that even if the SP order is known, a regret of $\Omega(U\sqrt{TK})$ is unavoidable. Furthermore, even if both the SP order and user peaks are known, every algorithm incurs a regret of $\Omega(U\sqrt{T}+\sqrt{TK})$.
@@ 1.2 Related Work
Our research intersects the combinatorial and matching threads of the multi-armed bandits literature, as well as single-peaked preferences from social choice theory. We provide a review of each of these strands of literature below.
@@
The study of single-peaked preferences dates back to seminal work in social choice theory [5, 1]. From a computational perspective, several problems have been investigated, including recognizing whether a given preferences profile admits a single-peaked order and constructing such an order when it exists [3, 13]. While recognition and construction of single-peaked orders can be solved efficiently, the problem becomes computationally intractable when preferences are nearly but not perfectly single-peaked. Specifically, finding an order that minimizes various distance measures from single-peakedness has been shown to be NP-hard [16, 11, 14]. In Section 4, we contribute to this literature by developing a technique to extract single-peaked orders from approximately single-peaked cardinal preferences.
@@ 2 Problem Definition
In this section, we formally introduce our model and notation (Section 2.1), address the case of general preferences (Section 2.2), and focus on single-peaked preferences (Section 2.3).
@@ 2.1 Model
The goal of the learner is to repeatedly recommend arms to users so as to maximize their overall satisfaction. Interaction proceeds over $T$ rounds. In each round $t\in[T]$, the learner chooses a matching $\pi_{t}:[U]\to[K]$, namely a function assigning each user a single item. The expected value of the matching is the total expected satisfaction of the users with their assigned arms,
$V(\pi_{t};\Theta)=\sum_{u\in[U]}\theta_{u,\pi_{t}(u)}.$
However, not all matchings are feasible. The algorithm has a budget constraint on the selected arms: each selected arm (i.e., there is at least one user $u$ such that $\pi_{t}(u)=k$) has a cost of $c_{k}$, and the total cost of the selected arms in every round should remain below a known budget $B\in\mathbb{N}$. Formally, for a matching $\pi$, let $\text{Im}(\pi)=\{k\in[K]:\exists u\in[U]\text{ s.t. }\pi(u)=k\}$. Hence, a matching is feasible only if $\sum_{k\in\text{Im}(\pi_{t})}c_{k}\leq B$. We denote the set of all feasible matchings as $\Pi$. Without loss of generality, we assume that $c_{k}\leq B$ for all $k\in[K]$, since otherwise such arms could never be selected. As in the motivating example of Section 1, the platform pays the content creators (arms) per creation, so multiple users can be matched to the same arm without incurring additional cost.
@@ 2.2 Warm Up: General Preferences
Without structural assumptions on $\Theta$, achieving low regret with polynomial time algorithms is hopeless. In particular, computing $\operatorname*{arg\,max}_{\pi\in\Pi}V(\pi;\Theta)$ is computationally intractable. Furthermore, due to a reduction from the maximum coverage problem [17], it is NP-hard to approximate the offline problem within any factor better than (11/e)(1-1/e). Since an efficient online algorithm with sublinear regret would imply an efficient approximation algorithm to the offline problem, such an online algorithm cannot exist. Due to space limitations, we defer the full analysis to Appendix E.
The literature typically addresses such cases, where the offline problem is NP-complete and thus cannot serve as a comparable benchmark, using the notion of $\alpha$-regret [8, 28]. Although computationally feasible, achieving sublinear $\alpha$-regret still implies linear regret.
In the online setting, the primary obstacle is computational rather than statistical. Indeed, optimal regret bounds can be achieved by applying the CUCB algorithm [8], which maintains element-wise UCB estimates for the reward matrix and repeatedly computes the optimal matching on these estimates. However, this optimization is NP-hard under general preferences, making CUCB computationally impractical. This motivates our assumption of structural properties for the expected reward matrix $\Theta$, which enables efficient optimization without sacrificing regret guarantees.
@@ 2.3 Single-Peaked Preferences
In other words, the entries of every row are unimodal. The index $p(u;\Theta)$ is called the peak of user $u$, or simply $p(u)$ when $\Theta$ is clear from context; if multiple indices attain the maximum, we fix an arbitrary choice. More generally,
@@
Note that any PSP matrix is also SP, with the identity order serving as its SP order. In contexts where the relevant order matters, we shall write that $\Theta$ is SP w.r.t. $\prec$. Furthermore, we say that a $\mathsf{CBR}$ instance is (P)SP if its expected reward matrix $\Theta$ is (P)SP.
@@
Figure 1: Illustration of PSP and SP instances. Each subfigure consists of curves representing the expected rewards of user-arm pairs. The instance in Figure 1(a) is SP, since if we reorder the arms, we get the PSP instance in Figure 1(b) .
To illustrate the PSP and SP properties, consider Figure 1. The instance comprises three users and five arms, and each subfigure illustrates the expected rewards for each user. The instance in Figure 1(a) is not PSP, since there are “valleys” that contradict (1). If we reorder the arms according to the order $\prec$ such that $1\prec 2\prec 5\prec 4\prec 3$, we obtain Figure 1(b). Indeed, we observe a unimodal preference shape for each user, indicating that the matrix is PSP. Thus, the instance in Figure 1(a) is SP and $\prec$ is its SP order.
We note that the SP order need not be unique; multiple SP orders may exist for the same matrix. For example, the reverse order of any SP order is also an SP order.
@@ 3 Offline Algorithm for SP instances
In this section, we present an efficient and optimal algorithm for computing $\operatorname*{arg\,max}_{\pi\in\Pi}V(\pi;\Theta)$ for SP instances. W.l.o.g., we shall assume that the SP order is the identity order, i.e., the instance is PSP. Indeed, an SP order can be extracted using our Extract-Order algorithm (see Algorithm 1) with $\varepsilon=0$ in $O(UK^{2})$ time, after which the columns can be reordered accordingly.
The main observation used by our algorithm relates to the optimal matching conditioned on a given subset of arms. Namely, we show that if the set of selected arms is $S$, the optimal matching assigns each user to the arm in $S$ that is closest to their peak. We formalize this in Lemma 1.
@@
Lemma 1 guides the design of a dynamic programming-based approach for finding the optimal matching for a given PSP matrix $P$, costs $c(\cdot)$, and budget $B$. We call this algorithm SP-Matching and explain its essence here, deferring full details to Appendix A. The algorithm maintains a table $F(k,b)$ representing the maximum achievable reward when arm $k$ is the rightmost selected arm, the total budget is at most $b$, and we consider only users $u$ such that $p(u)\leq k$. We add an auxiliary degenerate arm 0 such that $c_{0}=0$ and $P_{u,0}=0$ for every user $u$, and initialize $F(0,b)=0$ for every $b\in\{0\}\cup[B]$.
The recurrence relation for $F(k,b)$ follows from Lemma 1. When $k$ is the arm with the highest index among the selected arms and $i$ is the second highest, users with peaks before $i$ are unaffected by the inclusion of arm $k$, while those with peaks between $i$ and $k$ are assigned to whichever of the two arms offers a higher reward. Consequently, $F(k,b)$ is given by
@@
This maximum is always well-defined since arm 0 always satisfies its constraint. The optimal value is then
$V^{\star}=\max_{k\in[K]}\left[F(k,B)+\sum_{u:p(u)>k}P_{u,k}\right],$
where users with peaks beyond all selected arms are assigned to the last selected arm $k$. To enable efficient computation, we precompute for all pairs $(i,k)$ the sums $\sum_{u:i<p(u)\leq k}\max\{P_{u,i},P_{u,k}\}$ in $O(K^{2}U)$ time. This allows each entry in $F$ to be filled in $O(K)$ time, and since $F$ has $KB$ entries, the total runtime is $O(K^{2}U+K^{2}B)$.
@@
For any PSP matrix $P$, SP-Matching computes an optimal matching in time $O(K^{2}U+K^{2}B)$.
4 Online Algorithm for SP instances
@@
To tackle this challenge, we take a three-step approach. First, we define the concept of approximate single-peaked matrices that relaxes the strict SP condition while remaining amenable to analysis. Second, we present an efficient procedure to extract a plausible SP order from empirical data. Finally, we combine these tools in an explore-then-commit algorithm that achieves sublinear regret.

### v2

@@ 1.1 Our Contributions
We study a stochastic combinatorial bandit problem over $T$ rounds with $K$ arms and $U$ users. Each arm $k$ has a cost $c_{k}$, and the learner may select any subset of arms within a budget $B$, then assign users to the selected arms. Each user derives a stochastic satisfaction from each arm, and the goal is to maximize total user satisfaction. As the general case is NP-complete, we focus on SP preferences (see LABEL:{def:single-peaked-def}), which enable efficient solutions. Our main contributions are as follows.
SP instances are statistically challenging
To demonstrate that SP preferences do not trivialize the learning problem, we show that the worst-case regret remains unchanged compared to general preferences. Specifically, we prove that even when the SP order is known, a regret of $\Omega(U\sqrt{TK})$, which matches the lower bound for general preferences, is unavoidable. Furthermore, even if the user peaks are known as well, every algorithm incurs a regret of $\Omega(\max\{U\sqrt{T},\sqrt{TK}\})$.
@@
We consider the case of known SP order and user peaks (but not the cardinal preferences). Such an assumption is justified if, for example, preferences rely on a known ideological spectrum where user peaks correspond to stated or inferred inclinations. We design an efficient UCB-based algorithm, termed MvM, that achieves $\tilde{O}(U\sqrt{TK})$ regret. The key property used in our approach is the existence of a maximal preferences matrix in the confidence set of all plausible SP matrices—namely, a matrix that element-wise dominates all others in the set. We use this maximal matrix as the basis for optimistic matching using our offline algorithm.
@@ 1.2 Related Work
Our research intersects the combinatorial and matching threads of the multi-armed bandits literature, as well as single-peaked preferences from social choice theory.
@@
The study of single-peaked preferences dates back to seminal work in social choice theory [6, 1]. From a computational perspective, several problems have been investigated, including recognizing whether a given preferences profile admits a single-peaked order and constructing such an order when it exists [4, 14]. While recognition and construction of single-peaked orders can be solved efficiently, the problem becomes computationally intractable when preferences are nearly but not perfectly single-peaked. Specifically, finding an order that minimizes various distance measures from single-peakedness has been shown to be NP-hard [17, 12, 15]. In Appendix 6, we contribute to this literature by developing a technique to extract single-peaked orders from approximately single-peaked cardinal preferences.
@@ 2 Problem Definition
In this section, we formally introduce our model and notation (Section˜2.1), address the case of general preferences (Section˜2.2), and focus on single-peaked preferences (Section˜2.3).
@@ 2.1 Model
The goal of the learner is to recommend arms to users so as to maximize their overall satisfaction. Interaction proceeds over $T$ rounds. In each round $t\in[T]$, the learner chooses a matching $\pi_{t}:[U]\to[K]$, namely a function assigning each user a single item. The expected value of the matching is the total expected satisfaction of the users with their assigned arms, $V(\pi_{t};\Theta)=\sum_{u\in[U]}\theta_{u,\pi_{t}(u)}$.
However, not all matchings are feasible. The algorithm has a budget constraint on the selected arms: each selected arm (i.e., there is at least one user $u$ such that $\pi_{t}(u)=k$) has a cost of $c_{k}$, and the total cost of the selected arms in every round should remain below a known budget $B\in\mathbb{N}$. Formally, for a matching $\pi$, let $\text{Im}(\pi)=\{k\in[K]:\exists u\in[U]\text{ s.t. }\pi(u)=k\}$. Hence, a matching is feasible only if $\sum_{k\in\text{Im}(\pi_{t})}c_{k}\leq B$. We denote the set of all feasible matchings as $\Pi$. Without loss of generality, we assume that $c_{k}\leq B$ for all $k\in[K]$, since otherwise such arms could never be selected. As in the motivating example of Section˜1, the platform pays the content creators (arms) per creation, so multiple users can be matched to the same arm without incurring additional cost.
@@ 2.2 Warm Up: General Preferences
Without structural assumptions on $\Theta$, achieving low regret with polynomial time algorithms is hopeless. In particular, we establish the following hardness result based on a reduction from the maximum coverage problem [18]:
Theorem 1.
It is NP-hard to approximate $\operatorname*{arg\,max}_{\pi\in\Pi}V(\pi;\Theta)$ within any factor better than (11/e)(1-1/e).
Consequently, since an efficient online algorithm with sublinear regret would imply an efficient approximation algorithm for the offline problem, such an online algorithm cannot exist. We defer the full analysis to Appendix A. If we disregard computational constraints, optimal regret bounds can be attained by, e.g., applying the CUCB algorithm [9], assuming access to an exact optimization oracle. Formally,
Corollary 2.
There exists a UCB-based algorithm that achieves a regret of $O(U\sqrt{KT\log{T}})$, which is optimal up to logarithmic factors.
While this algorithm is statistically optimal, it assumes an optimization oracle that must solve an NP-hard problem at every time step, rendering it computationally impractical.
The literature typically addresses cases where the offline benchmark is intractable using the notion of $\alpha$-regret [9, 29, 30]. This metric uses the best efficiently computable approximation as the baseline, rather than the true optimum. Although it is computationally feasible to achieve sublinear $\alpha$-regret in our setting (e.g., with $\alpha=0.5$ using greedy oracle), this is a weaker guarantee compared to the standard regret. This limitation motivates our assumption of structural properties for $\Theta$, which enables efficient optimization without sacrificing regret guarantees.
@@ 2.3 Single-Peaked Preferences
In other words, the entries of every row are unimodal. The index $p(u;\Theta)$ is called the peak of user $u$, or simply $p(u)$ when $\Theta$ is clear from context; if multiple indices attain the maximum, we fix an arbitrary choice. [^1] See Appendix G for a discussion of how our algorithms and analyses handle such ties. More generally,
@@
Note that any PSP matrix is also SP, with the identity order serving as its SP order. In contexts where the relevant order matters, we shall write that $\Theta$ is SP w.r.t. $\prec$. Furthermore, we say that a $\mathsf{CBR}$ instance is (P)SP if its expected reward matrix $\Theta$ is (P)SP. The SP order need not be unique; for example, the reverse order of any SP order is also an SP order.
@@
Figure 1: Illustration of PSP and SP instances. Each subfigure consists of curves representing the expected rewards of user-arm pairs. The instance in Figure˜1(a) is SP, since if we reorder the arms, we get the PSP instance in Figure˜1(b) .
To illustrate the PSP and SP properties, consider Figure˜1. The instance comprises three users and five arms, and each subfigure illustrates the expected rewards for each user. The instance in Figure˜1(a) is not PSP, since there are “valleys” that contradict ˜(1). If we reorder the arms according to the order $\prec$ such that $1\prec 2\prec 5\prec 4\prec 3$, we obtain Figure˜1(b). Indeed, we observe a unimodal preference shape for each user, indicating that the matrix is PSP. Thus, the instance in Figure˜1(a) is SP and $\prec$ is its SP order.
3 Statistical Hardness of Single-Peaked Instances
While SP instances are highly structured, they remain statistically as challenging to learn. To demonstrate this, we establish a regret lower bound for SP instances of $\Omega(U\sqrt{TK})$, which matches the guarantee of Corollary˜2 for general preferences. Moreover, we provide a comparable lower bound for a more lenient case, where both the SP order and the users’ peaks are known to the learner in advance. We do so by constructing families of hard instances where the known structural properties are fixed, and only the cardinal values differ. Since the structural information is identical across instances, the learner cannot leverage it to distinguish between them. We present these lower bounds in Theorem˜3.
Theorem 3.
Theorem˜3 indicates that the SP assumption does not simplify the statistical aspect of the learning problem. However, as we demonstrate in the subsequent sections, it does affect the computational aspect–it enables efficient learning algorithms that achieve sublinear regret, standing in sharp contrast to the intractability of the general case discussed in Section˜2.2.
@@ 4 Offline Algorithm for SP instances
In this section, we present an efficient and optimal algorithm for computing $\operatorname*{arg\,max}_{\pi\in\Pi}V(\pi;\Theta)$ for SP instances. W.l.o.g., we shall assume that the SP order is the identity order, i.e., the instance is PSP. Indeed, an SP order can be extracted using our Extract-Order algorithm (see Algorithm˜2) with $\varepsilon=0$ in $O(UK^{2})$ time, after which the columns can be reordered accordingly.
The main observation used by our algorithm relates to the optimal matching conditioned on a given subset of arms. Namely, we show that if the set of selected arms is $S$, the optimal matching assigns each user to the arm in $S$ that is closest to their peak. We formalize this in Lemma˜4.
@@
Lemma˜4 guides the design of a dynamic programming-based approach for finding the optimal matching for a given PSP matrix $P$, costs $c(\cdot)$, and budget $B$. We call this algorithm SP-Matching and explain its essence here, deferring full details to Appendix C. The algorithm maintains a table $F(k,b)$ representing the maximum achievable reward when arm $k$ is the rightmost selected arm, the total budget is at most $b$, and we consider only users $u$ such that $p(u)\leq k$. We add an auxiliary degenerate arm 0 such that $c_{0}=0$ and $P_{u,0}=0$ for every user $u$, and initialize $F(0,b)=0$ for every $b\in\{0\}\cup[B]$.
The recurrence relation for $F(k,b)$ follows from Lemma˜4. When $k$ is the arm with the highest index among the selected arms and $i$ is the second highest, users with peaks before $i$ are unaffected by the inclusion of arm $k$, while those with peaks between $i$ and $k$ are assigned to whichever of the two arms offers a higher reward. Consequently, $F(k,b)$ is given by
@@
This maximum is always well-defined since arm 0 always satisfies its constraint. The optimal value is then $V^{\star}=\max_{k\in[K]}\left[F(k,B)+\sum_{u:p(u)>k}P_{u,k}\right]$, where users with peaks beyond all selected arms are assigned to the last selected arm $k$. To enable efficient computation, we precompute for all pairs $(i,k)$ the sums $\sum_{u:i<p(u)\leq k}\max\{P_{u,i},P_{u,k}\}$ in $O(K^{2}U)$ time. This allows each entry in $F$ to be filled in $O(K)$ time, and since $F$ has $KB$ entries, the total runtime is $O(K^{2}U+K^{2}B)$.
@@
For any PSP matrix, SP-Matching finds an optimal matching in time $O(K^{2}(U+B))$.
5 Online Algorithm for the Known SP Structure Regime
@@
In this section, we take a first step towards resolving this challenge by leveraging additional structural information of the problem. Specifically, we assume that both the SP order and user peaks are known in advance to the learner. Recall from Section˜3 that despite this knowledge, the learning problem remains statistically non-trivial. These structural assumptions, however, enable an efficient, optimistic UCB-style algorithm that achieves $\tilde{O}(U\sqrt{TK})$ regret. Our approach involves the construction of optimistic reward estimates using confidence sets [35, Chapter 8.4], as detailed in Section˜5.1, while preserving the single-peaked structure. Then, we use those estimates to design the algorithm described in Section˜5.2.
5.1 Confidence Sets and Maximal Matrix
Given the known SP order $\prec$ and user peaks $p(\cdot)$, we construct confidence sets for the expected reward matrix that contain the true parameters with high probability. Intuitively, these sets capture all matrices that are statistically plausible given the data observed so far, and naturally shrink as more data is collected. We begin by establishing notation for standard confidence bounds. Given a history $H_{t}$ at round $t$, we denote by $\bar{\theta}_{u,k}(t)$ and $n_{u,k}(t)$ the empirical mean reward and the number of pulls of the pair $(u,k)$ observed up to round $t$, respectively. The upper confidence bound is defined as $\textnormal{UCB}_{u,k}(t)=\bar{\theta}_{u,k}(t)+\sqrt{\nicefrac{{2\ln T}}{{n_{u,k}(t)}}}$, with the lower confidence bound $\textnormal{LCB}_{u,k}(t)$ defined analogously using subtraction. We formally define the confidence set as follows.
Definition 3 (Confidence set).
The confidence set $\mathcal{C}^{t}=\mathcal{C}^{t}(\prec,p,H_{t})$ contains a matrix $P$ if and only if it is consistent with $\prec$ and $p(\cdot)$, and for any $u,k$ it holds that $\textnormal{LCB}_{u,k}(t)\leq P_{u,k}\leq\textnormal{UCB}_{u,k}(t)$.
The standard UCB approach requires picking the optimistic matching at every time $t$, namely
This discrete optimization problem over a convex set may be tricky, as different matrices yield different optimal matchings. However, a structural property of $\mathcal{C}^{t}$ simplifies this optimization problem. As we demonstrate shortly in Lemma˜6, $\mathcal{C}^{t}$ includes a unique maximal matrix $P^{t}$ in an element-wise sense; thus, to find the solution of Equation˜2, we should only find the optimal matching w.r.t. $P^{t}$.
Lemma 6.
For any confidence set $\mathcal{C}^{t}(\prec,p,H_{t})$, there exists a unique element-wise maximal matrix $P^{t}\in\mathcal{C}^{t}$ such that $P^{t}_{u,k}\geq P_{u,k}$ for all $P\in\mathcal{C}^{t}$ and all $u\in[U],k\in[K]$. Furthermore, this matrix is given by $P^{t}_{u,k}=\begin{cases}\min_{i:k\preceq i\preceq p(u)}\textnormal{UCB}_{u,i}(t),&k\preceq p(u)\\ \min_{i:p(u)\preceq i\preceq k}\textnormal{UCB}_{u,i}(t),&p(u)\preceq k\end{cases}$.
We refer to the above matrix as the maximal matrix of $\mathcal{C}^{t}$, and note that its structure relies on knowledge of both the order $\prec$ and the peaks $p(\cdot)$.
5.2 The Match-via-Maximal Algorithm
Algorithm 1 Match-via-Maximal (MvM)
1:Order $\prec$, peaks $p^{\star}(\cdot)$
2:for $t=1$ to $T$ do
3:  Construct the maximal matrix $P^{t}$
[…]
5:  Observe rewards and update history $H_{t}$
Next, we present a combinatorial UCB-based algorithm, which we refer to as MvM (Algorithm˜1). MvM acts optimistically–in each round, it selects the optimal matching w.r.t. the corresponding maximal matrix. The regret analysis of MvM involves standard UCB-like arguments. By the construction of the confidence sets and applying Hoeffding’s inequality with a union bound over all $(u,k)$ pairs and time steps, the actual expected reward matrix $\Theta$ is in $\mathcal{C}^{t}$ for all $t\in[T]$ with probability at least $1-\frac{2KU}{T^{3}}$. Employing the clean event technique, we ignore the complementary (bad) event, which occurs with low probability. In every round $t$, we have
$\max_{\pi\in\Pi}V(\pi;\Theta)-V(\pi_{t};\Theta)\leq V(\pi_{t};P^{t})-V(\pi_{t};\Theta)\leq\sum_{u\in[U]}2\sqrt{\nicefrac{{2\ln T}}{{n_{u,\pi_{t}(u)}}}}.$ (3)
Summing Inequality ((3)) over all rounds, we obtain the following theorem:
Theorem 7.
For any SP instance with known SP order $\prec$ and peaks $p(\cdot)$, MvM achieves regret of at most $O(U\sqrt{TK\ln T})$ with per-round runtime of $O(K^{2}U+K^{2}B)$.
Extension to partial structural knowledge
[…]
6 Online Algorithm for the Unknown SP Structure Regime
We now turn to the more general setting, where the learner knows only that $\Theta$ is SP, but lacks knowledge of the specific underlying order or user peaks. When the order is unknown, the confidence set must include all matrices consistent with any valid SP order, and the optimistic approach from Section˜5 can no longer be applied efficiently; we discuss this barrier formally in Section˜6.4.
To tackle this challenge, we take a three-step approach. First, we define the concept of approximate single-peaked matrices that relaxes the strict SP condition while remaining amenable to analysis. Second, we present an efficient procedure to extract a plausible SP order from empirical data. Finally, we combine these tools in an explore-then-commit algorithm that achieves sublinear regret. A key insight underlying our approach is that due to estimation noise, we cannot hope to recover the exact SP order of the true matrix $\Theta$; instead, we aim to find some order under which the empirical estimates are approximately single-peaked, which suffices for near-optimal matching.